OpenStudy (anonymous):

18. A 2.0-gram bullet is shot into a tree stump. It enters at a speed of 3.00 x 104 cm/s and comes to rest after having penetrated 0.05 m in to the stump. What was the average force during the impact?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

I don't know. Sorry :/ maybe these users might be able to help you :) @abb0t @Abhisar @Michele_Laino @ParthKohli @iambatman

OpenStudy (anonymous):

now i must play the waiting game, thanks for you the suggestions.

TheSmartOne (thesmartone):

Haha, np :)

Parth (parthkohli):

\[\rm W = \Delta KE = F_{avg} \times distance\]

Parth (parthkohli):

I tried to put it in simple mathematical symbols. Do tell if you need further help.

OpenStudy (anonymous):

I'm in a pretty deep hole with this. I'm not sure where the numbers are plugged in.

OpenStudy (anonymous):

I would have to assume that: W = work; KE = kinetic energy; F = force but what does the triangle mean?

Parth (parthkohli):

OK, let's first start with the meaning of those symbols. In stopping the bullet, the tree is doing some work through reactionary force, friction and stuff. When work is done, there is always change in energy. Is it possible to track the change in that energy? It definitely is. The bullet possesses some kinetic energy as it goes into the tree. As it begins to penetrate, it loses the kinetic energy (remember - kinetic energy is the energy of motion). What exactly, in numerical terms, is this change? \(\)We can represent change in KE as \(\\\rm \Delta KE\), where \(\Delta\) (delta) is a symbol standing for "change". It is also a fact that work done through a distance is always force applied times that distance (displacement, actually). And this displacement has been given as 0.05 metres.

Parth (parthkohli):

\[\rm \Delta KE \equiv KE_{final} - KE_{initial}\]

OpenStudy (rhystic):

parthkohli is correct and because the bullet goes to rest the Vfinal is 0 meaning you have only the initial kinetic energy and you are just solving for force after.

OpenStudy (anonymous):

so KE final = 0

OpenStudy (anonymous):

how do i solve for the initial KE?

OpenStudy (rhystic):

KE = .5*m*v^2 you have all everything you need from the problem itself.

OpenStudy (anonymous):

I'm still confused, I used up 2 pages in my notebook trying to figure this out. What am i doing wrong?

OpenStudy (anonymous):

KEinitial =1/2*0.002kg*300^2m/s is this right?

Parth (parthkohli):

Yes, that is correct.

OpenStudy (anonymous):

but how do i simplify this?

OpenStudy (anonymous):

or does it need to be simplified at all?

OpenStudy (anonymous):

thats how it gets swaggin...

OpenStudy (anonymous):

\[DeltaKE = 0 - 1/2*0.002kg*300^2m/s\]

OpenStudy (rhystic):

you simplify it by multiplying the value out and getting a single number then use that back in the work equation to get the average force.

OpenStudy (anonymous):

how do i multiply them if there are no like terms? Or does that mathematical rule even apply in physics?

OpenStudy (rhystic):

what do you mean? -.5*.002*300^2 = what? the units combine to make kg*m^2/s^2 which is the unit for energy if that is what you are talking about.

OpenStudy (rhystic):

after you find that answer use that in W = deltaKE = F * distance .... you now have deltaKE and the distance is stated in the problem as .05m so you have F = (.5*.002*300^2)/.05

OpenStudy (anonymous):

(.5*.002*300^2)/5 = (.5 * .002*90000)/5 = 90/5 = 18

OpenStudy (anonymous):

so F = 18

OpenStudy (rhystic):

I didn't do the calculation but provided you inputed the numbers correctly then yes F = 18 N

OpenStudy (anonymous):

so you're saying there's a chance that the problem wasn't done correctly?

OpenStudy (rhystic):

no it was done incorrectly because you divide by 5m instead of .05 m the answer should be 1800 N

OpenStudy (rhystic):

all your conversions and everything else was right you just had the wrong distance in the last part.

OpenStudy (anonymous):

Thank You so much, this problem was giving me a rough day.